what are the steps to solve 3y over 2 plus y over 6 equals 35 over 3? It has been awhile since I've worked problems like this and I'm not forsure how to do it?
Answer by lwsshak3(11628)

what are the steps to solve 3y over 2 plus y over 6 equals 35 over 3? It has been awhile since I've worked problems like this and I'm not forsure how to do it?
**
3y/2+y/6=35/3
LCD:6
multiply every term by LCD
9y+y=70
10y=70
divide both sides by10
y=7